# 3INFN, Laboratori Nazionali del Gran Sasso, 67100 Assergi, Italy 4Max-Planck-Institut für Astronomie, Königstuhl 17, 69117, Heidelberg, Germany 5Universität Heidelberg, Zentrum für Astronomie (ZAH), Institut für Theoretische Astrophysik, Albert-Ueberle-Str. 2, 69120, Heidelberg, Germany

### v1

@@ 3.1 The metallicity threshold for (P)PISNe
Figure 4: Final helium core mass as a function of the ZAMS mass. The red lines are our new mesa tracks (models S23), while the blue lines correspond to results from parsec tracks (models C15). Dark (Light) gray shaded areas: regime fo PISNe (PPISNe). From top left to bottom right we show metallicities $Z=0.0001$, 0.0004, 0.0008, 0.001, 0.004, 0.008.
@@
In the S23 model (Fig. 2), even stars whose initial mass is higher than 250 M explode as PISNe if their metallicity lies between 0.001 and 0.002. At lower metallicities, stars with ZAMS mass $>250$ M collapse directly to black holes, because photodisintegration prevents the explosion, while stars with ZAMS mass between 110 and 230 M undergo PISNe. Stars with 64 $64\text{ M}_{\odot}\leq M_{\mathrm{ZAMS}}\leq 110\text{ M}_{\odot}$ undergo PPISN for $Z\leq 0.004$ and this mass range extends up to $M_{\mathrm{ZAMS}}=500\text{ M}_{\odot}$ for $0.002\leq Z\leq 0.004$.
@@
Figure 5: PISN (red lines) and PPISN (blue lines) rate densities as a function of redshift. The solid and dashed lines are obtained using the model by S23 and C15, respectively. The green shaded area represents the observational constraints set by Schulze et al. (2024). The black arrow highlights the upper limit from Schulze et al. (2024).
@@
Figure 6: PISN (upper panels) and PPISN (lower panels) rate density evolution as a function of redshift. The left-hand and right-hand plots show the results obtained with the new wind model by S23 and the one by C15, respectively. The black thick line represents the total rate density, whereas the colored lines show the contribution of individual metallicities from $Z=1\times 10^{-4}$ (violet solid line) to $Z=2\times 10^{-2}$ (dashed yellow line).
@@
At high redshift, the main contribution to the PISN rate comes from very metal-poor stars ($Z<0.002$), for which the mass-loss rates adopted in the S23 and C15 models behave almost in the same way (Fig. 6). However, the (P)PISN rates obtained with models S23 and C15 show some differences even at high redshift. Specifically, at redshift $z\geq{}8$, the rate of the S23 model becomes slightly higher than the rate of the C15 models. This happens because of other differences between the mesa and parsec stellar tracks that we already discussed in Section 3.1. Specifically, Fig. 4 shows that the final He core mass of the parsec tracks (C15) at metallicity $0.00060.0001\leq{}Z\leq{}$ 0.0006 tends to be lower than the one of the mesa tracks (Sabhahit et al. 2023), reducing the overall PISN rate in the C15 models.
@@ 3.3 The impact of metallicity evolution
Figure 6 shows the contribution of various metallicities to the (P)PISN rate for our two models. Below redshift $z=5$, the two models behave in a completely different way: if we assume the winds by Sabhahit et al. (2023) the main contribution to PISNe (PPISNe) comes from metallicity $Z=0.001-0.002$ ($Z=0.002$), whereas if we adopt the winds by Chen et al. (2015) the dominant metallicity is $Z=0.01$ for both PISNe and PPISNe. At $z>5$, the two models evolve in a more similar way and also lower metallicities become important.
Figure 6 assumes our fiducial metallicity spread $\sigma_{\mathrm{Z}}=0.4$ (see Eq. 7). In Figure 7, we show the impact of the parameter $\sigma_{\mathrm{Z}}$ on our results. Specifically, we change the metallicity spread $\sigma_{\mathrm{Z}}$ between 0.1 and 0.6.
@@
Figure 7: PISN (upper row) and PPISN (lower row) as a function of the redshift, varying the metallicity spread from $\sigma_{\mathrm{Z}}=0.1$ (dark blue dashed line) to $\sigma_{\mathrm{Z}}=0.6$ (yellow dashed line). Left-hand (right-hand) panels: models with stellar winds from S23 (from C15).
@@
$[\mathrm{yr}^{-1}\,\mathrm{Gyr}^{-3}]$ $[\mathrm{yr}^{-1}\,\mathrm{Gyr}^{-3}]$
@@
Table 1: PISN and PPISN rate density at redshift $z=0$ from B22, T23, and G24). For B22, models Empirical, eagle, IllustrisTNG and Millennium refer to different models for the star formation rate density evolution. For T23, models fiducial and $-3\sigma$ assume different rates of the 12C($\alpha$,$\gamma$)16O reaction. The different values reported for G24 assume three different values of the metallicity dispersion $\tilde{\sigma}_{\mathrm{Z}}$ and two different metallicity threshold for PISN $Z_{\text{th}}$.
@@ 4 Discussion
The stellar wind model proposed by S23 has a strong impact on the evolution of VMSs. The high mass-loss rate in the optically thick regime causes the most massive stars at metallicity $Z\geq{}0.004$ to end their life with a He core mass lower than 32 M, without entering the (pulsational) pair-instability regime. In contrast, at metallicity $Z=0.001$, even stars with masses between 250 and 500 M can undergo PISNe. This has important consequences on the expected mass function of black holes (Torniamenti et al., in prep.). Here, we have found that the enhanced mass-loss rate lowers the maximum metallicity for PISNe from $Z\sim 1.7\times 10^{-2}$ to $Z\sim 2\times 10^{-3}$, as also shown by S23. The absence of contributions by higher metallicities suppresses the predicted peak of the PISN rate at redshift $z\sim{2}$ and reduces the expected rate at redshift $z=0$ by two orders of magnitude compared to the parsec stellar tracks (C15). The PISN rate density of our mesa models falls within the observational constraints estimated by Schulze et al. (2024): 0.009 $0.009\text{ Gpc}^{-3}\mathrm{yr}^{-1}\leq\mathcal{R}_{\mathrm{PISN}}\leq 0.7\text{ Gpc}^{-3}\mathrm{yr}^{-1}$.
@@
To better comprehend the critical impact of stellar winds and metallicity on the (P)PISN rate, we compare our results with previous works. B22 demonstrated that the variation of the star formation rate model widely affects the PISN rate. Their empirical model aligns with our results from model C15, showing a peak around $z\sim 2$ and a similar value at $z=0$ of approximately $\mathcal{R}_{\mathrm{PISN}}\sim 6\text{ yr}^{-1}\,\mathrm{Gyr}^{-3}$. In the other models they considered the PISN rate peak is at $z\sim{3}$, resulting in a lower rate at $z=0$, but still exceeding the more recent observational constraints.
@@
T23 explored the impact of variations in the 12C($\alpha$,$\gamma$)16O reaction rate, predicting a peak of the PISN rate around $z\sim 6$. This higher redshift may be attributed to the different metallicity distribution they assumed. As already discussed by G24, binary interactions are not expected to significantly impact the predicted rate. Moreover, the inclusion of enhanced stellar winds can lead to wider binaries due to the large amount of mass ejected, reducing the effects of binary evolution.
@@
G24 also investigated how metallicity evolution influences the expected PISN rate. They assumed the metallicity distributed as a log-normal distribution around the fundamental metallicity relation (FMR; Curti et al. 2020) with a metallicity spread $\tilde{\sigma}_{\mathrm{Z}}$. Changing $\tilde{\sigma}_{\mathrm{Z}}$ from 0.15 to 0.7 and the maximum metallicity to have PISN $Z_{\text{th}}$ from $1.5\times 10^{-3}$ to $1.7\times 10^{-2}$, they observed a change of several orders of magnitude in the PISN rate, largely exceeding the upper limits by Schulze et al. (2024) when $\tilde{\sigma}_{\mathrm{Z}}\gtrsim{}0.35$ and $Z_{\text{th}}\geq 1.5\times 10^{-3}$. In our work, the metallicity spread $\sigma_{\mathrm{Z}}$ is defined as the standard deviation of a log-normal distribution around the average metallicity evolution of stars (Eq. 7). Hence, it is not directly comparable with the parameter $\tilde{\sigma}_{\mathrm{Z}}$ adopted by G24, even if both parameter quantify the dispersion of the stellar metallicity distribution (see the discussion by Sgalletta et al. 2024).
@@
A summary of all PISN rate densities at $z=0$ discussed here is provided in Table 1.
@@ 5 Summary
[…]
@@ Acknowledgements.
We wholeheartedly thank Gautham Sabhahit and collaborators for sharing their mesa inlists (https://github.com/Apophis-1/VMS_Paper1, https://github.com/Apophis-1/VMS_Paper2). Our work would not have been possible without this invaluable open-science input. We thank Marco Dall’Amico, Varsha Ramachandran, Andreas Sander, Raffaele Scala and Jorick Vink for useful discussions. MM, GI, and ST acknowledge financial support from the European Research Council for the ERC Consolidator grant DEMOBLACK, under contract no. 770017. MM and ST also acknowledge financial support from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) STRUCTURES. ST acknowledges financial support from the Alexander von Humboldt Foundation for the Humboldt Research Fellowship. GI acknowledges financial support from the La Caixa Foundation for the La Caixa Junior Leader fellowship 2024. GI also acknowledges financial support under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.4, - Call for tender No. 3138 of 18/12/2021 of Italian Ministry of University and Research funded by the European Union – NextGenerationEU. The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grants INST 35/1597-1 FUGG and INST 35/1503-1 FUGG. We use the mesa software (https://docs.mesastar.org/en/latest/) version r12115; (Paxton et al. 2011, 2013, 2015, 2018, 2019). We used sevn (https://gitlab.com/sevncodes/sevn) to generate our BBHs catalogs (Spera et al. 2019; Mapelli et al. 2020; Iorio et al. 2023). We used trackcruncher (https://gitlab.com/sevncodes/trackcruncher) (Iorio et al. 2023) to produce the tables needed for the interpolation in sevn. We used cosmorate (https://gitlab.com/Filippo.santoliquido/cosmo_rate_public) (Santoliquido et al. 2020, 2021) to compute the rate densities. This research made use of NumPy (Harris et al. 2020), SciPy (Virtanen et al. 2020), Pandas (The pandas development Team 2024). For the plots we used Matplotlib (Hunter 2007).
@@
Figure 10 shows the evolution of the He-core mass, while Table 2 reports the final values of the He core mass, at the end of core-He burning. At $Z=0.008$, the mass of the He core decreases quickly because of wind mass loss, and all models end their life with 13.7 $13.7\text{ M}_{\odot}\leq M_{\mathrm{He,f}}\leq 20\text{ M}_{\odot}$. These systems cannot enter the pair-instability regime and undergo a core-collapse supernova. Lowering the metallicity, at $Z=0.004$, only stars with 150 $150\text{ M}_{\odot}\leq M_{\text{ZAMS}}\leq 200\text{ M}_{\odot}$ can experience PPISNe, while in all the other cases we find $M_{\mathrm{He,f}}<32\text{ M}_{\odot}$. At $Z=0.001$, stars can experience different fates: only the $M_{\mathrm{ZAMS}}=50\text{ M}_{\odot}$ model undergoes a core-collapse supernova; models with $M_{\mathrm{ZAMS}}=75\,,\,100\text{ M}_{\odot}$ go through PPISN, while models with 125 $125\text{ M}_{\odot}\leq M_{\mathrm{ZAMS}}\leq 450\text{ M}_{\odot}$ can experience PISNe. Finally, if $M_{\mathrm{ZAMS}}=475\,,\,500\text{ M}_{\odot}$ the model directly collapses onto a black hole. At even lower metallicity ($Z=0.0008$), the initial stellar mass range of systems that can undergo a PISN is reduced: 125 $125\text{ M}_{\odot}\leq M_{\mathrm{ZAMS}}\leq 350\text{ M}_{\odot}$.
@@
For the winds of massive O-type stars, parsec adopts the fitting formulas by Vink et al. (2000) and Vink et al. (2001), with a correction introduced by Chen et al. (2015) to account for the effects of electron scattering, based on the models of Gräfener & Hamann (2008). In this framework, the mass-loss rate of an O-type star scales as $\dot{M}\propto Z^{\beta}$, where $Z$ is the metallicity and the exponent $\beta$ depends on the Eddington ratio $\Gamma_{\rm Edd}$: $\beta=0.85$ for $\Gamma_{\rm Edd}<2/3$, $\beta=2.45-2.4\times\Gamma_{\rm Edd}$ for $12/3\leq\Gamma_{\rm Edd}<$ 1, and $\beta=0.05$ for $\Gamma_{\rm Edd}\geq 1$ (Chen et al. 2015). This formalism provides a fit to the models by Gräfener & Hamann (2008) and Vink et al. (2011), incorporating the dependence of mass loss on the Eddington ratio.

### v2

@@ 3.1 The metallicity threshold for (P)PISNe
Figure 4: Final helium core mass as a function of the ZAMS mass. The orange lines are our new mesa tracks (models S23), while the blue lines correspond to results from parsec tracks (models C15). Dark (Light) gray shaded areas: regime for PISNe (PPISNe). From top left to bottom right we show metallicities $Z=0.0001$, 0.0004, 0.0008, 0.001, 0.004, 0.008.
@@
In the S23 model (Fig. 2), even stars whose initial mass is higher than 250 M explode as PISNe if their metallicity lies between 0.001 and 0.002. At lower metallicities, stars with ZAMS mass >250>250 M collapse directly to black holes, because photodisintegration prevents the explosion, while stars with ZAMS mass between 110 and 230 M undergo PISNe. Stars with $64\text{ M}_{\odot}\leq M_{\mathrm{ZAMS}}\leq 110\text{ M}_{\odot}$ undergo PPISN for $Z\leq 0.004$ and this mass range extends up to $M_{\mathrm{ZAMS}}=500\text{ M}_{\odot}$ for $0.002\leq Z\leq 0.004$.
@@
Figure 5: PISN (orange lines) and PPISN (blue lines) rate densities as a function of redshift. The solid and dashed lines are obtained using the model by S23 and C15, respectively. The green shaded area represents the observational constraints set by Schulze et al. (2024). The black arrow highlights the upper limit from Schulze et al. (2024).
@@
Figure 6: PISN (upper panels) and PPISN (lower panels) rate density evolution as a function of redshift. The left-hand and right-hand plots show the results obtained with the new wind model by S23 and the one by C15, respectively. The black thick line represents the total rate density, whereas the colored lines show the contribution of individual metallicities from $Z=1\times 10^{-4}$ (blue dashed line) to $Z=2\times 10^{-2}$ (solid yellow line).
@@
At high redshift, the main contribution to the PISN rate comes from very metal-poor stars ($Z<0.002$), for which the mass-loss rates adopted in the S23 and C15 models behave almost in the same way (Fig. 6). However, the (P)PISN rates obtained with models S23 and C15 show some differences even at high redshift. Specifically, at redshift $z\geq{}8$, the rate of the S23 model becomes slightly higher than the rate of the C15 models. This happens because of other differences between the mesa and parsec stellar tracks that we already discussed in Section 3.1. Specifically, Fig. 4 shows that the final He core mass of the parsec tracks (C15) at metallicity $0.0001\leq{}Z\leq{}0.0006$ tends to be lower than the one of the mesa tracks (Sabhahit et al. 2023), reducing the overall PISN rate in the C15 models.
@@ 3.3 The impact of metallicity evolution
Figure 6 shows the contribution of various metallicities to the (P)PISN rate for our two models. Below redshift $z=5$, the two models behave in a completely different way: if we assume the winds by Sabhahit et al. (2023) the main contribution to PISNe (PPISNe) comes from metallicity Z=0.0010.002Z=0.001-0.002 ($Z=0.002$), whereas if we adopt the winds by Chen et al. (2015) the dominant metallicity is $Z=0.01$ for both PISNe and PPISNe. At $z>5$, the two models evolve in a more similar way and also lower metallicities become important.
Figure 6 assumes our fiducial metallicity spread $\sigma_{\mathrm{Z}}=0.4$ (see Eq. 6). In Figure 7, we show the impact of the parameter $\sigma_{\mathrm{Z}}$ on our results. Specifically, we change the metallicity spread $\sigma_{\mathrm{Z}}$ between 0.1 and 0.6. Our fiducial value of $\sigma_{\mathrm{Z}}=0.4$ is compatible with observational calibrations (see, e.g., Chruslinska et al. 2019; Santoliquido et al. 2021). However, in Figure 7 we treat $\sigma_{\rm Z}$ as a parameter, given the large uncertainties and observational biases (see, e.g., Sgalletta et al. 2025, for a discussion).
@@
Gabrielli et al. (2024) found a similar behavior. They modeled the metallicity as a log-normal distribution centered on the fundamental metallicity relation (FMR; Curti et al. 2020), with a dispersion $\tilde{\sigma}_{\mathrm{Z}}$. When adopting the metallicity threshold $Z_{\text{th}}=0.017$, varying $\tilde{\sigma}_{\text{Z}}$ between 0.15 and 0.70 leads to nearly identical PISN rates at redshift $z=0$. In contrast, for a lower threshold ($Z_{\text{th}}=0.0015$), the same variation in $\tilde{\sigma}_{\text{Z}}$ results in a discrepancy of about five orders of magnitude (see Tab. 1). This indicates that the PISN rate is determined not only by $\sigma_{\text{Z}}$, but by the interplay between $\sigma_{\text{Z}}$ and the metallicity threshold $Z_{\text{th}}$.
Figure 7: PISN (upper row) and PPISN (lower row) as a function of the redshift, varying the metallicity spread from $\sigma_{\mathrm{Z}}=0.1$ (black dashed line) to $\sigma_{\mathrm{Z}}=0.6$ (yellow dashed line). Left-hand (right-hand) panels: models with stellar winds from S23 (from C15).
@@
$\mathrm{yr}^{-1}\,\mathrm{Gyr}^{-3}] [$$\mathrm{yr}^{-1}\,\mathrm{Gyr}^{-3}]$
@@
Table 1: PISN and PPISN rate density at redshift $z=0$ from B22, T23, and G24. For B22, models Empirical, eagle, IllustrisTNG and Millennium refer to different models for the star formation rate density evolution. For T23, models fiducial and $3\sigma$ assume different rates of the 12C($\alpha$$\gamma$)16O reaction. The different values reported for G24 assume three different values of the metallicity dispersion $\tilde{\sigma}_{\mathrm{Z}}$ and two different metallicity threshold for PISN $Z_{\text{th}}$.
@@ 4 Discussion
The stellar wind model proposed by S23 has a strong impact on the evolution of VMSs. The high mass-loss rate in the optically thick regime causes the most massive stars at metallicity $Z\geq{}0.004$ to end their life with a He core mass lower than 32 M, without entering the (pulsational) pair-instability regime. In contrast, at metallicity $Z=0.001$, even stars with masses between 250 and 500 M can undergo PISNe. This has important consequences on the expected mass function of black holes (Torniamenti et al., in prep.). Here, we have found that the enhanced mass-loss rate lowers the maximum metallicity for PISNe from $Z\sim 1.7\times 10^{-2}$ to $Z\sim 2\times 10^{-3}$, as also shown by S23. The absence of contributions by higher metallicities suppresses the predicted peak of the PISN rate at redshift $z\sim{2}$ and reduces the expected rate at redshift $z=0$ by two orders of magnitude compared to the parsec stellar tracks (C15). The PISN rate density of our mesa models falls within the observational constraints estimated by Schulze et al. (2024): $0.009\text{ Gpc}^{-3}\mathrm{yr}^{-1}\leq\mathcal{R}_{\mathrm{PISN}}\leq 0.7\text{ Gpc}^{-3}\mathrm{yr}^{-1}$.
@@
To better comprehend the critical impact of stellar winds and metallicity on the (P)PISN rate, we compare our results with previous works. A summary of all PISN rate densities at $z=0$ discussed here is provided in Table 1. B22 demonstrated that the variation of the star formation rate model widely affects the PISN rate. Their empirical model aligns with our results from model C15, showing a peak around $z\sim 2$ and a similar value at $z=0$ of approximately $\mathcal{R}_{\mathrm{PISN}}\sim 6\text{ yr}^{-1}\,\mathrm{Gyr}^{-3}$. In the other models they considered the PISN rate peak is at $z\sim{3}$, resulting in a lower rate at $z=0$, but still exceeding the more recent observational constraints.
@@
T23 explored the impact of variations in the 12C($\alpha$$\gamma$)16O reaction rate, predicting a peak of the PISN rate around $z\sim 6$. This higher redshift may be attributed to the different metallicity distribution they assumed. As already discussed by G24, binary interactions are not expected to significantly impact the predicted rate. Moreover, the inclusion of enhanced stellar winds can lead to wider binaries due to the large amount of mass ejected, reducing the effects of binary evolution.
@@
G24 also investigated how metallicity evolution influences the expected PISN rate. They assumed the metallicity distributed as a log-normal distribution around the fundamental metallicity relation (FMR; Curti et al. 2020) with a metallicity spread $\tilde{\sigma}_{\mathrm{Z}}$. Changing $\tilde{\sigma}_{\mathrm{Z}}$ from 0.15 to 0.7 and the maximum metallicity to have PISN $Z_{\text{th}}$ from $1.5\times 10^{-3}$ to $1.7\times 10^{-2}$, they observed a change of several orders of magnitude in the PISN rate, largely exceeding the upper limits by Schulze et al. (2024) when $\tilde{\sigma}_{\mathrm{Z}}\gtrsim{}0.35$ and $Z_{\text{th}}\geq 1.5\times 10^{-3}$. In our work, the metallicity spread $\sigma_{\mathrm{Z}}$ is defined as the standard deviation of a log-normal distribution around the average metallicity evolution of stars (Eq. 6). Hence, it is not directly comparable with the parameter $\tilde{\sigma}_{\mathrm{Z}}$ adopted by G24, even if both parameter quantify the dispersion of the stellar metallicity distribution (see the discussion by Sgalletta et al. 2025).
@@
[…]
In a companion paper (Boco et al. 2025), the authors argue that the S23 model, coupled with fast rotation or higher overshooting parameter, can also explain the existence of companion-less WR stars in the Small Magellanic Cloud (Schootemeijer et al. 2021). Additional tests and comparisons are needed, to address whether such models are compatible with other observational benchmarks, e.g. the Humphreys-Davidson limit (Humphreys & Davidson 1979).
@@ 5 Summary
In our S23 models, the PISN rate density grows up to a few $\times{}$ Gpc-3 yr-1 at redshift z=26z=2-6. The main contribution to PISNe and PPISNe comes from stars with metallicity Z=0.0010.002Z=0.001-0.002.
@@ Acknowledgements.
We wholeheartedly thank Gautham Sabhahit and collaborators for sharing their mesa inlists (https://github.com/Apophis-1/VMS_Paper1, https://github.com/Apophis-1/VMS_Paper2). Our work would not have been possible without this invaluable open-science input. We thank Guglielmo Costa, Marco Dall’Amico, Varsha Ramachandran, Andreas Sander, Raffaele Scala, Jorick Vink, Gautham Sabhahit, Daniel Pauli and Amedeo Romagnolo for useful discussions. MM, GI, and ST acknowledge financial support from the European Research Council for the ERC Consolidator grant DEMOBLACK, under contract no. 770017. MM and ST also acknowledge financial support from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) STRUCTURES. ST acknowledges financial support from the Alexander von Humboldt Foundation for the Humboldt Research Fellowship. GI acknowledges financial support from the La Caixa Foundation for the La Caixa Junior Leader fellowship 2024. GI also acknowledges financial support under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.4, - Call for tender No. 3138 of 18/12/2021 of Italian Ministry of University and Research funded by the European Union – NextGenerationEU. The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grants INST 35/1597-1 FUGG and INST 35/1503-1 FUGG. We use the mesa software (https://docs.mesastar.org/en/latest/) version r12115; (Paxton et al. 2011, 2013, 2015, 2018, 2019). We used sevn (https://gitlab.com/sevncodes/sevn) to generate our BBHs catalogs (Spera et al. 2019; Mapelli et al. 2020; Iorio et al. 2023). We used trackcruncher (https://gitlab.com/sevncodes/trackcruncher) (Iorio et al. 2023) to produce the tables needed for the interpolation in sevn. We used cosmorate (https://gitlab.com/Filippo.santoliquido/cosmo_rate_public) (Santoliquido et al. 2020, 2021) to compute the rate densities. This research made use of NumPy (Harris et al. 2020), SciPy (Virtanen et al. 2020), Pandas (The pandas development Team 2024). For the plots we used Matplotlib (Hunter 2007).
@@
Figure 10 shows the evolution of the He-core mass, while Table 2 reports the final values of the He core mass, at the end of core-He burning. At $Z=0.008$, the mass of the He core decreases quickly because of wind mass loss, and all models end their life with $13.7\text{ M}_{\odot}\leq M_{\mathrm{He,f}}\leq 20\text{ M}_{\odot}$. These systems cannot enter the pair-instability regime and undergo a core-collapse supernova. Lowering the metallicity, at $Z=0.004$, only stars with $150\text{ M}_{\odot}\leq M_{\text{ZAMS}}\leq 200\text{ M}_{\odot}$ can experience PPISNe, while in all the other cases we find $M_{\mathrm{He,f}}<32\text{ M}_{\odot}$. At $Z=0.001$, stars can experience different fates: only the $M_{\mathrm{ZAMS}}=50\text{ M}_{\odot}$ model undergoes a core-collapse supernova; models with $M_{\mathrm{ZAMS}}=75\,,\,100\text{ M}_{\odot}$ go through PPISN, while models with $125\text{ M}_{\odot}\leq M_{\mathrm{ZAMS}}\leq 450\text{ M}_{\odot}$ can experience PISNe. Finally, if $M_{\mathrm{ZAMS}}=475\,,\,500\text{ M}_{\odot}$ the model directly collapses onto a black hole. At even lower metallicity ($Z=0.0008$), the initial stellar mass range of systems that can undergo a PISN is reduced: $125\text{ M}_{\odot}\leq M_{\mathrm{ZAMS}}\leq 350\text{ M}_{\odot}$.
@@
For the winds of massive O-type stars, parsec adopts the fitting formulas by Vink et al. (2000) and Vink et al. (2001), with a correction introduced by Chen et al. (2015) to account for the effects of electron scattering, based on the models of Gräfener & Hamann (2008). In this framework, the mass-loss rate of an O-type star scales as $\dot{M}\propto Z^{\beta}$, where $Z$ is the metallicity and the exponent $\beta$ depends on the Eddington ratio $\Gamma_{\rm Edd}$: $\beta=0.85$ for $\Gamma_{\rm Edd}<2/3$, $\beta=2.45-2.4\times\Gamma_{\rm Edd}$ for $2/3\leq\Gamma_{\rm Edd}<1$, and $\beta=0.05$ for $\Gamma_{\rm Edd}\geq 1$ (Chen et al. 2015). This formalism provides a fit to the models by Gräfener & Hamann (2008) and Vink et al. (2011), incorporating the dependence of mass loss on the Eddington ratio.